Answer the question on the basis of the following information: Number of Total Marginal Workers Product Product 0 8 0 8 10 2 3 4 5 6 25 30 3 34 Refer to the data. The marginal product of the fourth worker:
Answer #1
No. Of Workers Total Product Marginal Product
0 0 -
1 8 8
2 18 10
3 25 7
4 30 5
5 33 3
6 34 1

The marginal product can be calculated as follows

large MP_L = rac{ riangle TP}{ riangle L}

MP-T

For 2 units of Labour

8 1 2 0

Thus, at L= 2, Total product = 18

At L = 3

MPL = (25 - 18)/(3-2) = 7/1 = 7

At L = 4

MPL = (30-25)/(4-3) = 5/1 = 5

At L = 5

MPL = (TP2 - 30)/(5-4)

3 = TP2 - 30

TP2 = 33

At L = 6

MPL = (34-33)/(6-5) = 1

Please contact if having any query thank you.